# Draw with Thought: Unleashing Multimodal Reasoning for Scientific Diagram Generation

### v1

@@ Abstract
Scientific diagrams are vital tools for communicating structured knowledge across disciplines. However, they are often published as static raster images, losing symbolic semantics and limiting reuse. While Multimodal Large Language Models (MLLMs) offer a pathway to bridging vision and structure, existing methods lack semantic control and structural interpretability, especially on complex diagrams. We propose Draw with Thought (DwT), a training-free framework that guides MLLMs to reconstruct diagrams into editable mxGraph XML code through cognitively-grounded Chain-of-Thought reasoning. DwT enables interpretable and controllable outputs without model fine-tuning by dividing the task into two stages: Coarse-to-Fine Planning, which handles perceptual structuring and semantic specification, and Structure-Aware Code Generation, enhanced by format-guided refinement. To support evaluation, we release Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations. Extensive experiments across eight MLLMs show that our approach yields high-fidelity, semantically aligned, and structurally valid reconstructions, with human evaluations confirming strong alignment in both accuracy and visual aesthetics, offering a scalable solution for converting static visuals into executable representations and advancing machine understanding of scientific graphics.
Figure 1: Overview of our Draw with Thought framework.
1 Introduction
@@
Scientific diagrams, such as model architectures, system workflows, and algorithmic flowcharts, are foundational to scholarly communication, providing compact visual representations of structured ideas [25]. However, most diagrams are published as rasterized PDF or PNG images [61], discarding their underlying symbolic structure and making them non-editable, non-executable, and difficult to interpret or reuse programmatically [12, 17, 52].
@@
Figure 2: Overview of our Draw with Thought. The framework for scientific diagram generation processes input diagrams through Coarse-to-Fine Planning (perceptual structuring and semantic layout planning) followed by Structure-Aware Code Generation to produce editable, high-fidelity diagram representations.
@@
Recovering structured representations from such diagrams entails mapping raw visual inputs to symbolic layouts comprising nodes, edges, spatial relations, and logical groupings [9]—a task that lies at the intersection of visual perception and structured abstraction. While Vision-Language Models (VLMs) have shown strong capabilities in aligning images with structured outputs, as demonstrated in captioning [31, 14], visual QA [47], and diagram layout prediction [43], their application to scientific diagram parsing remains nascent and largely limited to visually regular or low-complexity inputs. Existing studies have adopted various target formats that differ in expressiveness and abstraction level, such as SVG for vector rendering [44, 61], TikZ for LaTeX-integrated graphics [8, 9], and Python-based code for chart generation [55]. While SVG enables resolution-independent rendering, its reliance on low-level geometric primitives limits semantic interpretability [9]; TikZ supports precise layout control but introduces parsing challenges due to procedural syntax and structural variability [8, 9]; and although Python formats are effective for domain-specific plots, they lack the modularity and compositional flexibility needed to capture the heterogeneity of general-purpose scientific diagrams [55]. In contrast, XML-based formats such as mxGraph offer an explicit and editable graph abstraction, encoding diagrams as structured layouts with semantically grounded nodes, edges, and geometric constraints. This makes them particularly well-suited as target representations for recovering the symbolic structure of scientific diagrams from raw visual inputs [43].
@@
In summary, existing methods for scientific diagram parsing face three key limitations: (1) reliance on shallow or domain-specific formats that restrict structural expressiveness [61, 44]; (2) modeling paradigms that focus on layout detection or retrieval, lacking symbolic reasoning and structural supervision [61]; and (3) assumptions of clean vector inputs or templates [55], which hinder generalization to real-world, heterogeneous, rasterized diagrams.
Figure 3: Qualitative Comparison Between DwT and State-of-the-Art MLLMs. We visualize the output results. We observe that our model can understand the layout and the content of the image well, achieving better performance.
@@
To address these limitations, we present Draw with Thought, a training-free framework that leverages Multimodal Large Language Models (MLLMs) for symbolic reconstruction of scientific diagrams. Our approach draws inspiration from cognitive load theory [41], which suggests that breaking down complex tasks into manageable steps facilitates reasoning, and from structure mapping theory [22], which models analogy-making as a process of aligning relational structures. Building on such insights, we design a cognitively grounded Chain-of-Thought (CoT) prompting strategy that guides MLLMs through two key stages: Coarse-to-Fine Planning (Section 4.1), which includes Perceptual Structuring and Semantic Specification, and Structure-Aware Code Generation (Section 4.2), which incorporates iterative Format-Guided Refinement. These stages reflect how humans process diagrams—by sequentially recognizing visual elements, inferring their relationships, and assembling them into structured representations. Unlike prior works that rely on vision-only perception [55] or heuristic-based templates [16], our framework introduces a unified, interpretable reasoning pipeline compatible with diverse diagram styles and semantic structures. It supports high-fidelity generation of editable mxGraph XML code without any model fine-tuning. Our key contributions include:
@@
  • We introduce Draw with Thought, which guides MLLMs to generate executable mxGraph representations from rasterized scientific diagrams via cognitively grounded chain-of-thought prompting: covering visual decomposition, relational inference, and symbolic code synthesis.
@@
Table 1: Comparison of Scientific Diagram Datasets for Structural Understanding and Representation. We evaluate datasets across multiple dimensions. Full support ($\checkmark$) indicates comprehensive capability, partial support ($\checkmark$) indicates limited capability, and no support ($\times$) indicates absence of capability. Multimodality notation: T: Text-only, I: Image-only, I+T: Multi-modal with both images and text. Data authenticity reflects whether datasets contain real-world scientific diagrams from research publications and the complexity metric is expressed as the difficulty of data generation and evaluation.
@@
SciGraphQA[34] Scientific Charts $\times$ $\checkmark$ $\times$ $\times$ Text $\times$
Design2Code[46] Website UI $\checkmark$ $\times$ $\checkmark$ $\times$ HTML $\checkmark$
SketchFig[9] Hand-drawn Diagrams $\checkmark$ $\times$ $\checkmark$ $\times$ TikZ $\times$
MMCode[32] Algorithm Diagrams $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Code $\checkmark$
@@
2 Related Work
2.1 Image-to-Markup Conversion
Image-to-markup conversion supports a wide range of diagram types. Existing Studies rely on low-level vector primitives like SVG [44, 61], which ensure resolution independence but offer limited semantic detail. Others use Python-based chart-generation scripts including Plot2Code [55], PandasPlotBench [20] and Chart2Code [64], which excel in producing domain-specific plots yet struggle with heterogeneous diagrams [9]. Math-centric methods translate geometric figures or handwritten equations into  code [10, 53] and address symbolic inference, but often remain confined to specialized content. Besides, existing datasets largely fall short in covering key dimensions required for structured scientific diagram understanding. SciGraphQA [34] lacks visual inputs and complex structure, Design2Code [46] focuses on UI layouts with HTML rendering but lacks diagram-level abstraction, SketchFig [9] includes hand-drawn diagrams but provides limited layout semantics and MMCode [32] supports multimodal prompts but lacks authentic scientific content and XML-level structure. In contrast, XML-based structured markup language [43] under mxGraph standards [29], captures nodes, edges, and layout constraints in a unified form—an especially relevant feature for complex scientific diagrams that involve diverse graphical elements and rich relational structures. Consequently, to handle the intricacies of scientific diagram generation [25], our work adopts mxGraph to define the scientific diagram parsing task as an image-to-XML pipeline, enabling more expressive and editable representations.
2.2 Multimodal Large Language Models in Code Generation
Recent advances in Multimodal Large Language Models (MLLMs), such as Qwen-vl [6] and GPT-4 [28] demonstrate strong abilities to extract structured meaning from complex visual inputs, achieving notable results in image captioning [14, 31], visual question answering [47], and scene-level reasoning [21]. Meanwhile, code-oriented LLMs, including InCoder [19], StarCoder [33], and Code Llama [24], excel in generating executable programs across multiple languages. Existing studies that integrate visual cues with textual embeddings [61, 43] often focus on simpler diagrams or charts [55], where step-by-step reasoning involving Chain-of-Thought [54], self-refine [37, 16], self-reflect [13] reasoning-and-acting [62, 56]and plan-and-solve [51, 42] can enhance interpretability. However, many approaches rely on standardized domains [61] and struggle to capture the rich semantics of real-world scientific diagrams [9]. Finetuning models to generate lengthy code like XML [60], SVG [44, 61] or LaTeX [9] introduces substantial computational overhead, especially for small-parameter LLMs that must handle highly varied and domain-specific content. To overcome these challenges and facilitate XML-based code generation from complex scientific images, our work adopts a training-free strategy that leverages MLLMs for robust spatial reasoning, domain-specific logic, and knowledge-intensive tasks.
Inspired by cognitive load theory [41], which advocates breaking intricate problems into manageable steps-and structure mapping theory [22]—which models human reasoning through the alignment of relational structures—we propose Draw with Thought, a framework that guides MLLMs to parse and synthesize diagrammatic information without the need for expensive model finetuning.
3 Plot2XML: A Benchmark for Scientific Diagram-to-Code Generation
@@
3.1 Data Collection and Curation
We introduce Plot2XML, a novel benchmark dataset consisting of 247 scientific diagrams carefully selected from influential conference papers across multiple domains. The collection spans model architecture diagrams, flowcharts, system pipelines, and conceptual frameworks essential for scientific communication. Unlike SVG representations[15, 48], our Plot2XML diagrams in mxGraph format exhibit a geometric increase in both element count and code complexity, making this the first large-scale collection of scientific diagrams with manually verified markup-based structural annotations. We sampled diagrams from top-tier conferences, ensuring temporal and venue diversity that reflects contemporary visualization practices in research. Further details about the dataset construction and annotation process are provided in Appendix A.
3.2 Annotation Process
Each diagram was meticulously annotated with its corresponding mxGraph XML representation through a rigorous crowdsourcing process [50]. A team of expert annotators manually recreated each diagram using Draw.io (diagrams.net). The process involved analyzing the original raster image, identifying all visual elements and connections, recreating the diagram with high fidelity, exporting the structured XML representation, and verifying quality through secondary review. This manual recreation approach ensures perfect alignment between visual diagrams and their XML code representations, creating reliable image-code pairs critical for training and evaluating diagram understanding models.
3.3 Complexity Analysis
Figure 4: Complexity analysis of scientific diagrams. Visualized as a radar chart over five dimensions: Connection Complexity, Graphical Complexity, Color Effects, Text Annotation, and Special Elements. The chart compares three difficulty levels—Easy (blue, 85 diagrams, 34.4%), Medium (coral, 98 diagrams, 39.7%), and Hard (green, 64 diagrams, 25.9%)—highlighting distinct structural patterns across complexity axes
To systematically categorize the diagrams, we developed a multi-dimensional complexity analysis framework evaluating each diagram across five key dimensions: Connection Complexity (relationships between elements), Graphical Complexity (visual sophistication), Color Effects (semantic encoding), Text Annotation (textual density), and Special Elements (non-standard visual components). Based on these dimensions, we classified diagrams into three difficulty levels. As shown in Figure 4, Hard diagrams exhibit consistently higher complexity scores across all dimensions, with particularly pronounced differences in Connection Complexity ($score=4.1$) and Graphical Complexity ($score=4.3$). Further details about the dataset analysis are provided in Appendix A.
@@ 3.3 Complexity Analysis
5:$T_{percept}=\mathcal{M}_{MLLM}(D_{scientific},{Thought}_{percept})$
6:  $=(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector})$
7$\triangleright$ $T_{gestalt}$: perceptual grouping via Gestalt principles
8$\triangleright$ $T_{hierarchy}$: decomposition into visual primitives
9$\triangleright$ $T_{encoding}$: visual-to-semantic variable mapping
10$\triangleright$ $T_{connector}$: element linkage and routing topology
@@
12:$T_{hierarchy}=\mathcal{M}_{MLLM}(T_{percept},{Thought}_{hierarchy})$
@@
$\triangleright \mathcal{R}$: semantic regions (e.g., Input, Output)
$\triangleright \mathcal{E}$: typed elements with class labels (e.g., Entity)
$\triangleright \mathcal{L}$: layout constraints (alignment, layering, connectivity)
@@
19:$Y_{mxGraph}=\mathcal{M}_{MLLM}(T_{hierarchy},{Thought}_{\text{code}})$
20:  $=\bigcup_{i}Y_{i}$ =$(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{\text{edge}})$
21$\triangleright$ $Y_{\text{doc}}$: root declarations (mxfile, diagram)
22$\triangleright$ $Y_{\text{style}}$: visual style tokens (colors, borders, fonts)
23$\triangleright$ $Y_{\text{node}}$: nodes from $\mathcal{E}$ with geometry and IDs
24$\triangleright$ $Y_{\text{layout}}$: layout constraints from $\mathcal{L}$
25$\triangleright$ $Y_{\text{edge}}$: directed connectors with routing logic
@@
30:     if Draw.ioVerifier($Y_{mxGraph}^{(t)}$) == valid then break
@@
32$\triangleright$ Verification ensures XML schema validity, rendering success, and structural consistency
@@
34$\triangleright$ $T^{*}\leq T_{\text{refine}}$, earliest valid XML accepted by Draw.io
4 Draw with Thought
Following [51], We propose Draw with Thought (DwT), a planning-driven framework that transforms a rasterized scientific diagram $D_{scientific}$ into structured, editable graph code $Y_{mxGraph}$ in mxGraph XML format by guiding a Multimodal Large Language Model $\mathcal{M}_{MLLM}$ through planning-and-generate Chain of Thought. Inspired by how humans decompose diagrams through sequential reasoning [41, 22]—from perceptual recognition to semantic understanding and symbolic execution—we explicitly separate the diagram-to-code task into two interpretable stages  [^1] Please see Appendix C.1 for prompt details.: (1) Coarse-to-Fine Planning involving ${Thought}_{percept}$ and ${Thought}_{hierarchy}$ (Section 4.1). (2) Structure-Aware Generation involving ${Thought}_{code}$ and ${RefineThought}_{format}$ (Section 4.2), as detailed in Algorithm 1 and Figure 2. This separation enables the model to reason over symbolic abstractions before emitting graphical code.
4.1 Coarse-to-Fine Planning for Structured Visual Understanding

### v2

@@ Abstract.
Scientific diagrams are vital tools for communicating structured knowledge across disciplines. However, they are often published as static raster images, losing symbolic semantics and limiting reuse. While Multimodal Large Language Models (MLLMs) offer a pathway to bridging vision and structure, existing methods lack semantic control and structural interpretability, especially on complex diagrams. We propose Draw with Thought (DwT), a training-free framework that guides MLLMs to reconstruct diagrams into editable mxGraph XML code through cognitively inspired Chain-of-Thought reasoning. DwT enables interpretable and controllable outputs without model fine-tuning by dividing the task into two stages: Coarse-to-Fine Planning, which handles perceptual structuring and semantic specification, and Structure-Aware Code Generation, enhanced by format-guided refinement. To support evaluation, we release Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations. Extensive experiments across eight MLLMs show that our approach yields high-fidelity, semantically aligned, and structurally valid reconstructions, with human evaluations confirming strong alignment in both accuracy and visual aesthetics, offering a scalable solution for converting static visuals into structurally valid and renderable representations and advancing machine understanding of scientific graphics.
Multimodal Large Language Models, Chain-of-Thought, Code Generation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755731ccs: Computing methodologies Planning with abstraction and generalization
Figure 1. Overview of our Draw with Thought framework.
1. Introduction
@@
Scientific diagrams, such as model architectures, system workflows, and algorithmic flowcharts, are foundational to scholarly communication, providing compact visual representations of structured ideas (Hegarty et al., 1991). However, most diagrams are published as rasterized PDF or PNG images (Xing et al., 2024), discarding their underlying symbolic structure and making them non-editable, non-executable, and difficult to interpret or reuse programmatically (Chen et al., 2024; Favreau et al., 2016; Wang and Lian, 2021).
@@
Figure 2. Overview of our Draw with Thought. The framework for scientific diagram generation processes input diagrams through Coarse-to-Fine Planning (perceptual structuring and semantic layout planning) followed by Structure-Aware Code Generation to produce editable, high-fidelity diagram representations.
@@
Recovering structured representations from such diagrams entails mapping raw visual inputs to symbolic layouts comprising nodes, edges, spatial relations, and logical groupings (Belouadi et al., 2024)—a task that lies at the intersection of visual perception and structured abstraction. While Vision-Language Models (VLMs) have shown strong capabilities in aligning images with structured outputs, as demonstrated in captioning (Li et al., 2023; Dai et al., 2023), visual QA (Song et al., 2022), and diagram layout prediction (Roberts et al., 2024), their application to scientific diagram parsing remains nascent and largely limited to visually regular or low-complexity inputs. Existing studies have adopted various target formats that differ in expressiveness and abstraction level, such as SVG for vector rendering (Rodriguez et al., 2023; Xing et al., 2024), TikZ for LaTeX-integrated graphics (Belouadi et al., [n. d.], 2024), and Python-based code for chart generation (Wu et al., 2024a). While SVG enables resolution-independent rendering, its reliance on low-level geometric primitives limits semantic interpretability (Belouadi et al., 2024); TikZ supports precise layout control but introduces parsing challenges due to procedural syntax and structural variability (Belouadi et al., [n. d.], 2024); and although Python formats are effective for domain-specific plots, they lack the modularity and compositional flexibility needed to capture the heterogeneity of general-purpose scientific diagrams (Wu et al., 2024a). In contrast, XML-based formats such as mxGraph offer an explicit and editable graph abstraction, encoding diagrams as structured layouts with semantically grounded nodes, edges, and geometric constraints. This makes them particularly well-suited as target representations for recovering the symbolic structure of scientific diagrams from raw visual inputs (Roberts et al., 2024).
@@
In summary, existing methods for scientific diagram parsing face three key limitations: (1) reliance on shallow or domain-specific formats that restrict structural expressiveness (Xing et al., 2024; Rodriguez et al., 2023); (2) modeling paradigms that focus on layout detection or retrieval, lacking symbolic reasoning and structural supervision (Xing et al., 2024); and (3) assumptions of clean vector inputs or templates (Wu et al., 2024a), which hinder generalization to real-world, heterogeneous, rasterized diagrams.
@@
To address these limitations, we present Draw with Thought, a training-free framework that leverages Multimodal Large Language Models (MLLMs) for symbolic reconstruction of scientific diagrams. Our approach draws inspiration from cognitive load theory (Plass et al., 2010), which suggests that breaking down complex tasks into manageable steps facilitates reasoning, and from structure mapping theory (Gentner, 1983), which models analogy-making as a process of aligning relational structures. Building on such insights, we design a cognitively inspired Chain-of-Thought (CoT) prompting strategy that guides MLLMs through two key stages: Coarse-to-Fine Planning (Section 4.1), which includes Perceptual Structuring and Semantic Specification, and Structure-Aware Code Generation (Section 4.2), which incorporates iterative Format-Guided Refinement. These stages reflect how humans process diagrams—by sequentially recognizing visual elements, inferring their relationships, and assembling them into structured representations. Unlike prior works that rely on vision-only perception (Wu et al., 2024a) or heuristic-based templates (Fang et al., 2025), our framework introduces a unified, interpretable reasoning pipeline compatible with diverse diagram styles and semantic structures. It supports high-fidelity generation of editable mxGraph XML code without any model fine-tuning. Our key contributions include:
@@
  • We introduce Draw with Thought, which guides MLLMs to generate executable mxGraph representations from rasterized scientific diagrams via cognitively inspired chain-of-thought prompting: covering visual decomposition, relational inference, and symbolic code synthesis.
@@
2. Related Work
2.1. Image-to-Markup Conversion
Image-to-markup conversion supports a wide range of diagram types. Existing Studies rely on low-level vector primitives like SVG (Rodriguez et al., 2023; Xing et al., 2024), which ensure resolution independence but offer limited semantic detail. Others use Python-based chart-generation scripts including Plot2Code (Wu et al., 2024a), PandasPlotBench (Galimzyanov et al., 2024) and Chart2Code (Zhao et al., 2025), which excel in producing domain-specific plots yet struggle with heterogeneous diagrams (Belouadi et al., 2024). Math-centric methods translate geometric figures or handwritten equations into code (Blecher et al., [n. d.]; Wang and Liu, 2020) and address symbolic inference, but often remain confined to specialized content. Besides, existing datasets largely fall short in covering key dimensions required for structured scientific diagram understanding. SciGraphQA (Li and Tajbakhsh, 2023) lacks visual inputs and complex structure, Design2Code (Si et al., 2024) focuses on UI layouts with HTML rendering but lacks diagram-level abstraction, SketchFig (Belouadi et al., 2024) includes hand-drawn diagrams but provides limited layout semantics and MMCode (Li et al., 2024) supports multimodal prompts but lacks authentic scientific content and XML-level structure. In contrast, XML-based structured markup language (Roberts et al., 2024) under mxGraph standards (Kaplan and Rabelo, 2024), captures nodes, edges, and layout constraints in a unified form—an especially relevant feature for complex scientific diagrams that involve diverse graphical elements and rich relational structures. Consequently, to handle the intricacies of scientific diagram generation (Hegarty et al., 1991), our work adopts mxGraph to define the scientific diagram parsing task as an image-to-XML pipeline, enabling more expressive and editable representations.
2.2. Multimodal Large Language Models in Code Generation
Recent advances in Multimodal Large Language Models (MLLMs), such as Qwen-vl (Bai et al., 2025) and GPT-4 (Hurst et al., 2024) demonstrate strong abilities to extract structured meaning from complex visual inputs, achieving notable results in image captioning (Dai et al., 2023; Li et al., 2023), visual question answering (Song et al., 2022), and scene-level reasoning (Gao et al., 2024). Meanwhile, code-oriented LLMs, including InCoder (Fried et al., [n. d.]), StarCoder (Li et al., [n. d.]), and Code Llama (Grattafiori et al., 2023), excel in generating executable programs across multiple languages(Zhang et al., 2024; Yang et al., 2024). Existing studies that integrate visual cues with textual embeddings (Xing et al., 2024; Roberts et al., 2024) often focus on simpler diagrams or charts (Wu et al., 2024a), where step-by-step reasoning involving Chain-of-Thought (Wei et al., 2022), self-refine (Madaan et al., 2023; Fang et al., 2025), self-reflect (Cheng et al., 2024) reasoning-and-acting (Yao et al., [n. d.]; Wu et al., 2025)and plan-and-solve (Wang et al., 2023; Qiao et al., 2024) can enhance interpretability. However, many approaches rely on standardized domains (Xing et al., 2024) and struggle to capture the rich semantics of real-world scientific diagrams (Belouadi et al., 2024). Finetuning models to generate lengthy code like XML (Xie et al., 2024), SVG (Rodriguez et al., 2023; Xing et al., 2024) or LaTeX (Belouadi et al., 2024) introduces substantial computational overhead, especially for small-parameter LLMs that must handle highly varied and domain-specific content. To overcome these challenges and facilitate XML-based code generation from complex scientific images, our work adopts a training-free strategy that leverages MLLMs for robust spatial reasoning, domain-specific logic, and knowledge-intensive tasks.
Inspired by cognitive load theory (Plass et al., 2010), which advocates breaking intricate problems into manageable steps-and structure mapping theory (Gentner, 1983)—which models human reasoning through the alignment of relational structures—we propose Draw with Thought, a framework that guides MLLMs to parse and synthesize diagrammatic information without the need for expensive model finetuning.
Table 1. Comparison of Scientific Diagram Datasets for Structural Understanding and Representation. We evaluate datasets across multiple dimensions. Full support ($\checkmark$) indicates comprehensive capability, partial support ($\checkmark$) indicates limited capability, and no support ($\times$) indicates absence of capability. Multimodality notation: T: Text-only, I: Image-only, I+T: Multi-modal with both images and text. Data authenticity reflects whether datasets contain real-world scientific diagrams from research publications and the complexity metric is expressed as the difficulty of data generation and evaluation.
@@ 2.2. Multimodal Large Language Models in Code Generation
SciGraphQA(Li and Tajbakhsh, 2023) Scientific Charts $\times$ $\checkmark$ $\times$ $\times$ Text $\times$
Design2Code(Si et al., 2024) Website UI $\checkmark$ $\times$ $\checkmark$ $\times$ HTML $\checkmark$
SketchFig(Belouadi et al., 2024) Hand-drawn Diagrams $\checkmark$ $\times$ $\checkmark$ $\times$ TikZ $\times$
MMCode(Li et al., 2024) Algorithm Diagrams $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Code $\checkmark$
@@
3. Plot2XML: A Benchmark for Scientific Diagram-to-Code Generation
@@
3.1. Data Collection and Curation
We introduce Plot2XML, a novel benchmark dataset consisting of 247 scientific diagrams carefully selected from influential conference papers across multiple domains. The collection spans model architecture diagrams, flowcharts, system pipelines, and conceptual frameworks essential for scientific communication. Unlike SVG representations(Eisenberg and Bellamy-Royds, 2014; Song et al., 2025), our Plot2XML diagrams in mxGraph format exhibit a geometric increase in both element count and code complexity, making this the first large-scale collection of scientific diagrams with manually verified markup-based structural annotations.
3.2. Complexity Analysis
Figure 3. Complexity analysis of scientific diagrams. Visualized as a radar chart over five dimensions: Connection Complexity, Graphical Complexity, Color Effects, Text Annotation, and Special Elements. The chart compares three difficulty levels—Easy (blue, 85 diagrams, 34.4%), Medium (coral, 98 diagrams, 39.7%), and Hard (green, 64 diagrams, 25.9%)—highlighting distinct structural patterns across complexity axes.
To systematically categorize the diagrams, we developed a multi-dimensional complexity analysis framework evaluating each diagram across five key dimensions: Connection Complexity (relationships between elements), Graphical Complexity (visual sophistication), Color Effects (semantic encoding), Text Annotation (textual density), and Special Elements (non-standard visual components). Based on these dimensions, we classified diagrams into three difficulty levels. As shown in Figure 3, Hard diagrams exhibit consistently higher complexity scores across all dimensions, with particularly pronounced differences in Connection Complexity ($score=4.1$) and Graphical Complexity ($score=4.3$).
@@ 3.2. Complexity Analysis
[…]
6:  $(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector})$
7:  $\triangleright$ $T_{gestalt}$: perceptual grouping via Gestalt principles
8:  $\triangleright$ $T_{hierarchy}$: decomposition into visual primitives
9:  $\triangleright$ $T_{encoding}$: visual-to-semantic variable mapping
10:  $\triangleright$ $T_{connector}$: element linkage and routing topology
@@
[…]
@@
14:  $\triangleright$ $\mathcal{R}$: semantic regions (e.g., Input, Output)
15:  $\triangleright$ $\mathcal{E}$: typed elements with class labels (e.g., Entity)
16:  $\triangleright$ $\mathcal{L}$: layout constraints (alignment, layering, connectivity)
@@
[…]
20:  $\bigcup_{i}Y_{i} =$$(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{\text{edge}})$
21:  $\triangleright$ $Y_{\text{doc}}$: root declarations (mxfile, diagram)
22:  $\triangleright$ $Y_{\text{style}}$: visual style tokens (colors, borders, fonts)
23:  $\triangleright$ $Y_{\text{node}}$: nodes from $\mathcal{E}$ with geometry and IDs
24:  $\triangleright$ $Y_{\text{layout}}$: layout constraints from $\mathcal{L}$
25:  $\triangleright$ $Y_{\text{edge}}$: directed connectors with routing logic
@@
[…]
@@
[…]
@@
[…]
4. Draw with Thought
Following (Wang et al., 2023), We propose Draw with Thought (DwT), a planning-driven framework that transforms a rasterized scientific diagram $D_{scientific}$ into structured, editable graph code $Y_{mxGraph}$ in mxGraph XML format by guiding a Multimodal Large Language Model $\mathcal{M}_{MLLM}$ through planning-and-generate Chain of Thought. Inspired by how humans decompose diagrams through sequential reasoning (Plass et al., 2010; Gentner, 1983)—from perceptual recognition to semantic understanding and symbolic execution—we explicitly separate the diagram-to-code task into two interpretable stages: (1) Coarse-to-Fine Planning involving ${Thought}_{percept}$ and ${Thought}_{hierarchy}$ (Section 4.1). (2) Structure-Aware Generation involving ${Thought}_{code}$ and format (Section 4.2), as detailed in Algorithm 1 and Figure 2. This separation enables the model to reason over symbolic abstractions before emitting graphical code.
4.1. Coarse-to-Fine Planning for Structured Visual Understanding